OpenStudy (anonymous):

x+3=19-3x pls show working

OpenStudy (anonymous):

Bring everything on one side x + 3 - 19 + 3x = 0 Sum 4x - 16 = 0 Isolate x 4x = 16 Solve! :) x = 4

OpenStudy (anonymous):

4x=16 then divide by 4 x=4 then plug it back in 4+3=19-12 check